# GAI_arXiv.bib

### v2

@@ Abstract
This study examines whether collective reasoning among generative agents can facilitate novel and coherent thinking that leads to innovation. To achieve this, it proposes GAI, a new LLM-empowered framework designed for reflection and interaction among multiple generative agents to replicate the process of innovation. The core of the GAI framework lies in an architecture that dynamically processes the internal states of agents and a dialogue scheme specifically tailored to facilitate analogy-driven innovation. The framework’s functionality is evaluated using Dyson’s invention of the bladeless fan as a case study, assessing the extent to which the core ideas of the innovation can be replicated through a set of fictional technical documents. The experimental results demonstrate that models with internal states significantly outperformed those without, achieving higher average scores and lower variance. Notably, the model with five heterogeneous agents equipped with internal states successfully replicated the key ideas underlying the Dyson’s invention. This indicates that the internal state enables agents to refine their ideas, resulting in the construction and sharing of more coherent and comprehensive concepts.
@@
Keywords: generative AI, Large language models, Multi-Agent Systems, Innovation, Patent
I Introduction
@@
Generative agents, including LLM agents, can engage in reflections for decision-making or analysis by breaking down complex tasks and conducting systematic reasoning. LLM-based Multi-agent System (LLM-MAS) extends this capability by adding collective reasoning functions, with applications already emerging in fields such as software development and social simulations ([chan2023chatevalbetterllmbasedevaluators, guo2024largelanguagemodelbased, hong2024metagptmetaprogrammingmultiagent, park2023generativeagentsinteractivesimulacra, qian2024chatdevcommunicativeagentssoftware, sun2024llmbasedmultiagentreinforcementlearning, Xiong_2023, zhou2024sotopiainteractiveevaluationsocial]). One particularly noteworthy area is the application of LLM-MAS to scientific discussions. In some studies, generative agents have been used to address tasks including mathematical problems and medical debates, with interesting findings such as agents correcting their mistakes through dialogue ([tang2024medagentslargelanguagemodels, du2023improvingfactualityreasoninglanguage]). Applying these capabilities of generative agents to discussions based on technical documents, such as patents, could potentially replicate the collective interactions that human teams engage in during the innovation process.
@@
Efforts to analyze the innovation process have historically included observational experiments in actual corporate environments, but these have faced significant limitations due to low reproducibility and difficulties in manipulating experimental conditions. As a result, the effects of variables such as organizational structure, communication methods, and degree of diversity on the innovation process remain not fully understood. If LLM-MAS can replicate the innovation process, it would enable repeated simulations with adjustable conditions under high reproducibility.
@@
Furthermore, it is conceivable that artificial intelligence technologies, including LLMs, will increasingly be applied in real-world corporate innovation settings, creating a new domain of practice that could be described as “AI-driven innovation.” If LLM-MAS can simulate the innovation process, it would also be possible to pioneer this new domain by integrating human teams with generative agent teams, demonstrating new possibilities for AI-human collaboration in driving innovation.
@@
The objective of this study is to examine whether collective reasoning among generative agents can facilitate novel and coherent thinking that leads to innovation. To achieve this, the study specifically focuses on analogy-driven innovation and proposes GAI, a new LLM-empowered framework designed for reflection and interaction among multiple generative agents to replicate this process.
@@
The core of the GAI framework lies in an architecture that dynamically processes the internal states of agents and a dialogue scheme specifically designed to facilitate analogy-driven innovation. In the former, agents generate and refine ideas based on intrinsic motivation and critical introspection, leveraging previously reviewed documents and interactions with others to form their current thoughts, which serve as the basis for their contributions to the dialogue. In the latter, the scheme incorporates insights from management science and cognitive science to replicate the process of identifying functional similarities and mechanical differences, enabling the transfer of solutions across diverse technical domains.
@@
To evaluate the functionality of the framework, this study uses Dyson’s invention of the bladeless household fan as a case study. It examines the extent to which analogy-driven innovations can be replicated by the proposed framework using a set of fictional technical documents.
II Proposed framework
II-A Agent architecture
@@
The architecture of each agent consists of two primary modules driven by an LLM: a memory module and an internal state module (Figure 1). The memory module records the contents of reviewed technical documents, past conversations with others, and the agent’s prior reflections, all timestamped for reference. These records are then organized into a structured memory, which is input into the internal state module. The internal state module processes newly acquired technical information and statements from others during the dialogue, drawing on both recent thoughts and memories retrieved from the memory module. Based on this, the internal state module forms and updates the agent’s current thoughts, which serve as the foundation for its future reflections and statements.
@@
When forming and updating current thoughts, the internal state module undergoes two processes: idea generation and introspection.
During the idea generation process, the module generates multiple ideas for reflections on a technological domain based on the memories received from the memory module and the agent’s thoughts at that time. Subsequently, the internal state module evaluates these ideas on a scale of 1 to 10 for each of three criteria—Novelty, Importance, and Consensus—using a language model and selects ideas that maximize an intrinsic reward calculated as a weighted sum of these criteria:
@@
where $r_{ij}$ represents the intrinsic reward of agent $i$, $x_{ij}$ is its idea on object $j$, and $\alpha_{i}$, $\beta_{i}$, and $\gamma_{i}$ respectively represent the weights for the 1-10 evaluation on each criterion, which are denoted by Novelty$(x_{ij})$, Importance$(x_{ij})$, and Consensus$(x_{ij})$ in Equation (1). The Novelty criterion assesses the degree to which an idea introduces new content that has not been present in prior conversations and reflections. The Importance criterion evaluates how significant the idea is in light of previous discussions and reflections. The Consensus criterion measures the extent to which the idea aligns with the opinions of the majority of participants, based on past conversations. By assigning different weights to these criteria for each agent, the framework can represent diverse intrinsic motivations. For instance, an agent with a higher weight on Novelty will prioritize curiosity and innovation, whereas an agent with a higher weight on Consensus will emphasize collaboration and agreement.
During the introspection process, the generated idea are critically examined to identify any technical ambiguities or contradictions using a language model. The internal state module then adjusts the idea to address these issues, refining it to improve clarity and coherence.
@@
The internal state module not only reflects on the agent’s own thoughts about technological domains but also evaluates the opinions expressed by other participants. Specifically, based on prior conversations and the agent’s current thoughts, it identifies aspects of each participant’s opinions that the agent agrees with or finds interesting, as well as those that the agent considers contradictory or lacking.
@@
As a result, the agent’s current thoughts are continuously updated, incorporating refined ideas and reflections on others. The format of the current thoughts is given in Appendix -A. During the dialogue, the agent makes statements based on its current thoughts, shaped through these processes, as well as on memories retrieved from the memory module. These current thoughts are also used when reflecting on new information or statements acquired or heard during the dialogue. Even if agents are initially assigned homogeneous roles and personality settings, their thoughts dynamically evolve in distinct ways through these processes.
II-B Dialogue scheme for Analogy-driven innovation
@@
Innovation takes various forms, such as technology-driven innovation, market-driven innovation, open, or closed innovation. This paper focuses on Design by Analogy (DbA), an approach to innovation that has been extensively studied across disciplines, including management science and cognitive science ([holyoak1994, gassmann2008opening, linsey2008modality, fu2015design, fu2013meaning, mcadams2002quantitative, song2022design, song2019design, hope2017accelerating, jeong2014creating, moreno2014fundamental, murphy2014function, linsey2008increasing, goel1997design, huhns1988argo, Melnychuk2020]). Many groundbreaking innovations, such as the application of racing car technologies to running shoes and various forms of biomimicry, employ cross-domain analogies. DbA is a method that addresses problems in a target domain by extracting and transferring design solutions from other domains (source domains). The central focus of this approach lies in identifying similarities and differences. In particular, it is widely recognized that superior innovations abstract away from surface-level or attribute-based similarities, instead relying on functional and relational similarities for analogy ([holyoak1994, linsey2008modality, fu2015design, fu2013meaning, mcadams2002quantitative, bhatta1996design]).
@@
Specifically, this scheme facilitates the following stages in sequence:
Preparation: Reviewing technical documents from different domains. Each agent is provided with multiple technical documents belonging to the target domain and source domains and is tasked with reviewing them.
Phase A: Extracting functional similarities. Through multiple rounds of discussion, the agents identify functional similarities between the target domain and the source domains. They also determine which of these similarities are more central to the problem at hand.
Phase B: Extracting mechanical differences. Building on the functional similarities identified in Phase A, the agents examine the mechanisms that generate these shared functions in the target and source domains. Through several rounds of discussion, they identify the differences in these mechanisms and evaluate which of these differences are most significant.
Phase C: Transferring solutions. In this phase, the agents explore how the mechanical differences identified in Phase B can be applied to address the challenges in the target domain. Through multiple rounds of discussion, they determine what aspects of the source domains’ mechanisms need to be modified and incorporated into the target domain, as well as what features of the target domain’s existing functions or mechanisms should be retained or discarded.
Phase D: Identifying new challenges. As the agents deliberate on the solution transfer discussed in Phase C, they also assess whether any unforeseen challenges might arise during the application of these solutions.
Phase E: Identifying new opportunities. In contrast to Phase D, this phase focuses on exploring whether the solution transfer creates new value beyond solving the existing challenges in the target domain.
A sample of the dialogue prompts designed for facilitating this process is provided in Appendix -B.
II-C Organizational Structure
@@
III Experiment
In the experiment, we use Dyson’s invention of the bladeless household fan, Dyson Air Multiplier, as an example of innovation and evaluate the extent to which the proposed framework can replicate the key elements of this invention’s ideas. The evaluation is conducted in a sandbox environment using fictional technical documents.
III-A Case: Dyson Air Multiplier
@@
In household fans, it is common to cover the spinning blades with a grid-like guard to prevent accidental contact. However, such guards not only reduce the efficiency and consistency of airflow but also make the device less convenient to carry and clean. Dyson Air Multiplier addressed these challenges in household fans by applying the principles of fluid dynamics used in jet engines to household fans and achieving a transition from a physical blade-based airflow mechanism to a fluid dynamics-based airflow mechanism utilizing pressure differential.
@@
The Dyson Air Multiplier is often referred to as a ”bladeless fan,” but this label is somewhat misleading ([hallgeisler2022]). In reality, the device uses blades enclosed within the base, which are powered by an efficient brushless motor to draw in air. Atop the base, a ring-shaped airfoil with a diameter of approximately 30 cm is mounted. Air drawn into the device is expelled through the narrow slits of a jet nozzle on the inner surface of the ring. During this process, the fan pulls in air from behind the ring, effectively ”multiplying” the airflow ([mansson2014flow]). Specifically, the air drawn from the base is accelerated through the jet nozzle, creating a pressure gradient that induces surrounding air to be pulled in (inducement). Additionally, the high-speed expelled air entrains the surrounding air, generating a consistent airflow directed forward from the ring (entrainment).
@@
Dyson introduced the first generation of the Dyson Air Multiplier to the market in 2009. However, the first-generation device faced challenges with noise levels. To address this issue, Dyson invested over \$60 million in further research and development, culminating in the release of the second generation in 2014. The second-generation model incorporated Helmholtz cavities into the base, effectively resolving the turbulence issues that were a primary source of noise ([guardian2014, hallgeisler2022]).
III-B Experimental Method
@@
In the experiment, dialogues among generative agents are conducted with household fans as the target domain and industrial ejectors as the source domain. An ejector is an industrial device used for transporting fluids (liquids or gases). It is utilized not only for water supply, exhaust, and smoke extraction but also for neutralization, dilution, and cooling through the mixing of gases or liquids, as well as for creating vacuum conditions as a vacuum pump.
@@
Although the application domains are entirely different, household fans and industrial ejectors share a functional similarity in the movement of fluids. However, the mechanisms that enable this function are entirely distinct. In a typical ejector, a nozzle draws high-speed driving fluid into the body as the driving source. As the high-speed driving fluid reduces the pressure inside the body, creating negative pressure, the surrounding low-speed external fluid is pulled into the body. The mixed fluid of the driving and external fluids is then discharged outside the body through a diffuser. This allows the ejector to move external fluid without requiring any external power source.
@@
In the practical stage of the proposed framework, actual technical documents, such as patent filings, are expected to be used. However, in the sandbox environment presented in this paper, fictional technical documents based on real-world technical information are prepared. These fictional documents, like actual patent documents, include miscellaneous information that does not necessarily lead to the target invention. For instance, in the case of industrial ejectors, the document contains information about a wide range of industrial applications far removed from household use, details about unrelated components, and data on various types of irrelevant fluids and materials. Similarly, the technical documents for household fans include various unrelated information. Agents must extract fragments of information that lead to a ”solution” from this broad and miscellaneous data and connect those fragments using their independent reasoning. In this particular experiment, the agents must reinterpret the concept of driving fluid from the perspective of a household fan and devise a method to generate it within the constraints of household technology and energy environments. However, the agents are prohibited, by prompts, from using pre-trained knowledge about Dyson products or any information available on the internet. Instead, they are required to base their statements and reflections strictly on logical conclusions or associations derived from their prior discussions and reflections.
III-C Evaluations
@@
The evaluation is conducted for the solutions derived by the agents from the dialogue using a point-based system based on the nine criteria listed in Table I (maximum of 8 points). These criteria are developed from the perspective of whether the transfer of solutions from industrial ejectors to household fans comprehensively and coherently covers the technological elements corresponding to the first-generation Dyson Air Multiplier. Since the elements of the second generation may require computational fluid dynamics simulations for the specific shape of the fan to address the turbulence issues, they were excluded from the evaluation in this experiment, which only utilizes language models.
Criterion Description

### v3

@@ Abstract
This study examines whether collective reasoning among generative agents can facilitate novel and coherent thinking that leads to innovation. To achieve this, it proposes GAI (Generative Agents for Innovation), a new LLM-empowered framework designed for reflection and interaction among multiple generative agents to replicate the process of innovation. The core of the GAI framework lies in an architecture that dynamically processes the internal states of agents and a dialogue scheme specifically tailored to facilitate analogy-driven innovation. The framework’s functionality is evaluated using Dyson’s invention of the bladeless fan as a case study, assessing the extent to which the core ideas of the innovation can be replicated through agents’ dialogue using a set of fictional technical documents. The experimental results demonstrate that models with internal states significantly outperformed those without, achieving higher average scores and lower variance. Notably, the model with five heterogeneous agents equipped with internal states successfully replicated the key ideas underlying the Dyson’s invention. This indicates that the framework enables agents to refine their ideas, resulting in the construction and sharing of more coherent and comprehensive concepts.
{IEEEkeywords}
@@
Generative AI, Innovation, Large language models, Multi-Agent Systems, Patents
1 INTRODUCTION
@@
Generative agents, including LLM agents, can perform various complex tasks by breaking them down, conducting systematic reasoning, leveraging external resources, learning from experiences, and planning for future. LLM-based Multi-agent System (LLM-MAS) extends this capability by incorporating collective reasoning functions, with applications already emerging in fields such as software development, urban planning, scientific debates, and social simulations ([1, 7, 8, 9, 16, 12, 19, 32, 35, 41, 43, 44, 45, 46, 48, 49, 53, 65, 67, 68, 73, 76, 80, 86, 87, 88, 90, 91, 96, 95, 98, 103]). Applying these capabilities of generative agents to discussions based on technical documents, such as patents, could potentially replicate the collective interactions that human teams engage in during innovation processes.
@@
Efforts to analyze innovation processes have historically included observational experiments in actual corporate environments, but these have faced significant limitations due to low reproducibility and difficulties in manipulating experimental conditions. As a result, the effects of conditions such as diversity in team member, communication methods, and organizational structure on innovation processes remain not fully understood. If LLM-MAS can replicate the innovation processes of human teams, it would enable repeated simulations under adjustable conditions with high reproducibility.
@@
Furthermore, artificial intelligence technologies, including LLMs, will increasingly be applied in real-world corporate innovation settings ([2, 4, 5, 6, 13, 15, 22, 23, 29, 30, 31, 57, 69, 71, 74, 75, 81, 82, 94]). If LLM-MAS can simulate innovation processes, it would also be possible to pioneer this new domain by integrating generative agent teams with human teams, demonstrating new possibilities for AI-human collaboration in driving innovation.
@@
The objective of this study is to examine whether collective reasoning among generative agents can facilitate novel and coherent thinking that leads to innovation. To achieve this, the study focuses on analogy-driven innovation and proposes GAI (Generative Agents for Innovation), a new LLM-empowered framework designed for reflection and interaction among generative agents to replicate this process.
@@
The core of the GAI framework lies in an agent’s architecture that dynamically processes its internal state, and a dialogue scheme specifically designed to facilitate analogy-driven innovation. In the former, each agent generates and refines ideas based on intrinsic motivation and critical introspection, leveraging previously reviewed documents and interactions with others. This process forms and updates the current thought of the agent, allowing the agent’s contributions to the dialogue to dynamically self-evolve in a coherent way over time. In the latter, the scheme incorporates insights from management and cognitive science to replicate the process of identifying functional similarities and mechanical differences, enabling the transfer of solutions across various technological domains.
@@
To evaluate the functionality of the framework, this study performs two experiments using a set of fictional technical documents. The one is to take Dyson’s invention of the bladeless fan as a case study and examine the extent to which the innovation can be replicated using household fans as the target domain and industrial ejectors as the source domain. The other is to apply the framework to ten different source domains to examine its applicability.
2 PROPOSED FRAMEWORK
2.1 AGENT ARCHITECTURE
@@
The architecture of an agent consists of two primary LLM-driven modules: a memory module and an internal state module (Figure 1). The memory module records a) contents of reviewed technical documents, b) past conversations with others, and c) the agent’s prior reflections, all timestamped for reference. These records are then organized into structured memory, which is input into the internal state module.
The internal state module processes newly acquired technical information and statements from others during the dialogue and facilitates innovative and coherent thinking of an agent. Specifically, taking in external inputs from technical documents and other agents, as well as internal inputs from the agent’s prior reflections and its latest thoughts, the module continuously updates the agent’s current thoughts, which is a collection of texts structured according to the dialogue scheme for analogy-driven innovation. Current thoughts recurrently serve as the foundation for future reflections of the agent and function as the basis for the agent’s statement in a dialogue, providing it with strong context awareness and coherency.
@@
When forming and updating current thoughts, the internal state module undergoes two processes: idea generation and introspection. During the idea generation process, the module is prompted to generate multiple ideas distinct with each other for reflections on a technological domain based on the memories retrieved from the memory module and the agent’s current thoughts at that time. Subsequently, the module initiates another LLM instance and prompts it to evaluate these ideas by an integer score on a scale of 1 to 10 for each of three criteria:
  • Novelty: the degree to which the generated idea introduces new content that has not been present in prior conversations and reflections
  • Importance: the degree to which the generated idea is significant in light of previous discussions and reflections
  • Consensus: the degree to which the generated idea aligns with the opinions of the majority of participants, based on past conversations
The module selects the idea that maximizes an intrinsic reward calculated as a weighted sum of these criteria:
@@
[…]
During the introspection process, the selected idea is critically examined to identify any technical ambiguities or contradictions. Specifically, the module sets up another LLM instance as an internal critic and prompts it to examine if there are any important points in the selected idea that are unclear from the perspective of technology, and if there are any contradictions from the perspective of technology. The internal state module is then prompted to update the idea by including responses that accurately address the questions and points of contradiction raised through the introspection, refining the idea to improve clarity and coherence.
@@
The internal state module not only reflects on the agent’s own thoughts about technological domains but also evaluates the opinions expressed by other participants. Specifically, based on prior conversations and the agent’s current thoughts, the module is prompted to identify aspects of each participant’s opinions that the agent agrees with or finds interesting, as well as those that it considers contradictory or lacking. These reflections on others are then reflected in the agent’s current thoughts together with its own ideas.
@@
Through these steps, the agent’s current thoughts are continuously updated, incorporating refined ideas and reflections on domains and others. During the dialogue, the agent makes statements based on its current thoughts, shaped through these processes, as well as on memories retrieved from the memory module. Even if agents are initially assigned homogeneous roles and personality, their thoughts dynamically evolve in distinct ways through these processes.
2.2 DIALOGUE SCHEME FOR ANALOGY-DRIVEN INNOVATION
@@
Innovation takes various forms, such as technology-driven innovation, market-driven innovation, open or closed innovation. This paper focuses on Design by Analogy (DbA), an approach to innovation that has been extensively studied across disciplines, including management science and cognitive science ([17, 18, 20, 25, 42, 33, 37, 39, 50, 51, 58, 59, 61, 63, 77, 78]). Many groundbreaking innovations, such as the application of racing car technologies to running shoes and various forms of biomimicry, employ cross-domain analogies. DbA addresses problems in a target domain by extracting and transferring design solutions from other domains (source domains). The central focus of this approach lies in identifying similarities and differences. In particular, it is widely recognized that superior innovations abstract away from surface-level or attribute-based similarities, instead relying on functional and relational similarities for analogy ([3, 18, 17, 42, 51, 58]).
@@
Specifically, this scheme facilitates the procedure shown in Figure 2.
Figure 2: Dialogue scheme for analogy-driven innovation
2.3 ORGANIZATIONAL STRUCTURE
@@
3 EXPERIMENTS
3.1 CASE: Dyson Air Multiplier
@@
In household fans, it is common to cover the spinning blades with a grid-like guard to prevent accidental contact. However, such guards not only reduce the efficiency and consistency of airflow but also make the device less convenient to carry and clean. Dyson Air Multiplier addressed these challenges applying the principles of fluid dynamics used in jet engines and achieving a transition from a physical-blade-based airflow mechanism to a fluid-dynamics-based airflow mechanism leveraging pressure differential.
@@
Air Multiplier is often referred to as a ”bladeless fan,” but this is somewhat misleading ([28]). In reality, it uses blades enclosed within the base, powered by an efficient brushless motor to draw in air. Atop the base, a ring-shaped airfoil with a diameter of approximately 30 cm is mounted. Air drawn into the device is expelled through narrow slits of the jet nozzle on the inner surface of the ring. During this process, the fan pulls in air from behind the ring, effectively ”multiplying” the airflow ([56]). Specifically, the air drawn from the base is accelerated through the jet nozzle, creating a pressure gradient that induces surrounding air to be pulled in (inducement). Additionally, the high-speed expelled air entrains the surrounding air, generating a consistent airflow directed forward from the ring (entrainment).
@@
Dyson introduced the first generation of Air Multiplier to the market in 2009. However, it faced challenges with noise levels. To address this issue, Dyson invested over \$60 million in further R&D, culminating in the release of the second generation in 2014, which incorporated Helmholtz cavities into the base, effectively resolving the turbulence issues that were a primary source of noise ([26, 28]).
3.2 EXPERIMENT 1
@@
In the experiment, dialogues are conducted with household fans as the target domain and industrial ejectors as the source domain. Ejectors are industrial devices used for transporting fluids (liquids or gases). They are utilized not only for water supply, exhaust, and smoke extraction but also for neutralization, dilution, and cooling through the mixing of gases or liquids, as well as for creating vacuum conditions as a vacuum pump.
@@
Although the applications are entirely different, the tow domains share a functional similarity, the movement of fluids. However, the mechanisms that enable this function are entirely distinct. In a typical ejector, a nozzle draws high-speed driving fluid into the body as a driving source. As the high-speed driving fluid reduces the pressure inside the body, creating negative pressure, the surrounding low-speed external fluid is pulled into the body. The mixed fluid of the driving and external fluids is then discharged outside the body through a diffuser. This allows the ejector to move external fluid without any external power sources.
@@
In the practical stage of the proposed framework, actual technical documents, such as patent filings, are expected to be used. However, in the experimental environment presented in this paper, fictional technical documents based on real-world technical information are prepared, primarily considering budgetary constraints. These fictional documents, like actual patent documents, include miscellaneous information that does not necessarily lead to the target invention. For instance, the document for ejectors contains information about a wide range of industrial applications far removed from household use, details about unrelated components, and data on various types of irrelevant fluids and materials. Similarly, the document for household fans include various unrelated information. Agents must extract fragments of information that lead to a ”solution” from this broad and miscellaneous data and connect those fragments using their independent reasoning. In this particular experiment, to reach the target innovation, agents are required to reinterpret the concept of driving fluid from the perspective of a household fan and devise a method to generate it within the constraints of household technology and energy environments. However, the agents are prohibited, by prompts, from using pre-trained knowledge about Dyson products or any information available on the internet. Instead, they are required to base their statements and reflections strictly on logical conclusions or associations derived from their prior discussions and reflections.
@@
The evaluation is conducted for the solutions derived from the dialogue using a point-based system based on the nine criteria listed in Table 1 (maximum of 8 points). These criteria are developed from the perspective of whether the transfer of solutions comprehensively and coherently covers the technological elements corresponding to the first-generation Dyson Air Multiplier. Since the second generation requires computational fluid dynamics simulations for the specific shape of the fan to address the turbulence issues, they were excluded from the evaluation in this experiment, which only utilizes language models.
Table 1: Evaluation criteria
No. Description